A. What is the resistance of an electric frying pan that draws 12 amperes of current when connected to a 120 V circuit?



B. If your skin were very moist so that your resistance was only 1000 ohms, and you touched the terminals of a 24 V battery, how much current would you draw?



C. A certain battery has an 18.0 V EMF and an internal resistance of 0.015 Ω. Find the current, terminal voltage, and power dissipated by a load with 0.500 Ω.



with solution, thank you!



1
Expert's answer
2022-04-19T15:27:36-0400

a) According to the Ohm's law:


"V = IR"

where "R" is the resistance, "I = 12A" is the current and "V = 120V" is the voltage. Thus, obtain:


"R = \dfrac{V}{I} = 10\Omega"

b) Similarly:


"I = \dfrac{V}{R} = \dfrac{24V}{1000\Omega} = 0.024A"

c) The current in this case is (see https://courses.lumenlearning.com/physics/chapter/21-2-electromotive-force-terminal-voltage/):


"I = \dfrac{emf}{R_{load} +r}"

where "emf = 18V, R_{load}=0.5\Omega, r=0.015\Omega". Thus, obtain:


"I = \dfrac{18V}{0.5\Omega +0.015\Omega} \approx 35.0A"

The terminal voltage:


"V = emf-Ir = 18V-35A\cdot 0.015\Omega \approx 17.5V"

The power:


"P = I^2R_{load} = (35.0A)^2\cdot 0.5\Omega \approx 611W"